Chapter 1: Problem 1

  1. Suppose that real GDP is currently $97 billion per year and natural real GDP is currently $100 billion. Measured as a percentage, what is the GDP gap?
  2. Suppose natural real GDP is growing by $4 billion per year. By how much must real GDP have risen after two years to close the GDP gap?
Answer #1

A) GDP gap is the deviation of gdp from full employment level GDP or potential GDP or natural real GDp.

GDP gap=100-97=3 billion

GDP gap as percent=3/100*100=3%

B) natural GDP Increases by 4 billion each year ,so after two year natural GDP will increase by 8 billion.

And new natural GDP=100+8=108 billion.

New GDP gap=108-97=11 billion

So real GDP has rise by 11 billion to close the GDP.

So growth of real GDP in two years to close the GDP gap=11/97)*100=11.34%